# Boundary terms and on-shell action in Ricci-based gravity theories: the Hamiltonian formulation

### v1

@@ I Introduction
There is one issue that has remained unsolved over all these years though, namely, the development of a well-defined Hamiltonian formulation for any RBG. The main pitfall behind lied on formal aspects, and mainly on establishing a well-defined variational principle for these theories. The point is that as far as one assumes fixed values of the fields at the boundaries of the volume over which the action is integrated, the boundaries terms, as the Gibbons-Hawking-York (GHY) term of GR, are removed, since variations of the connection turn out null at the boundaries. Nevertheless, the presence of such boundary terms have been shown for the case of Palatini $f(R)$ gravities to be fundamental, among other stuff, to recover the correct expression for the Schwarzschild black hole entropy Saez-ChillonGomez:2020afj . In the same way, such boundary terms, analogs of the GHY term, have been obtained by different approaches BarberoG:2021cei ; Gomez:2021roj .
The aim of this paper is to go one step further in the enlightening of this topic with respect to previous analysis. Indeed, here we shall consider a completely general RBG theory and we settle to obtain first the corresponding boundary terms in the gravitational action and, second, to establish a well-defined Hamiltonian formalism for this class of theories. To do so, the mapping of RGB into GR-like form is used in order to transform the general gravitational action into one with just linear terms of the Ricci tensor Afonso:2018bpv . Then, by decomposing the free symmetric connection into two parts, one compatible with a metric and the other a 3-rank tensor, the Riemann and Ricci tensors are expressed in terms of these two connections for foliating the corresponding manifold into a family of hypersurfaces. The latter allows to express such tensors in terms of extrinsic and intrinsic magnitudes of the hypersurfaces, the so-called Gauss-Codazzi equations (also known as the $3+1$ decomposition), already implemented successfully in Palatini $f(R)$ gravities Gomez:2021roj . Finally, the Hamiltonian formulation is obtained by following the standard procedure of implementing the Arnowitt-Deser-Misner (ADM) decomposition. Results mimic GR formulation, as expected due to the mapping. However, important different features arise, as shown when computing the ADM energy for the Schwarzschild black hole.
@@ II Variational principle in Ricci-based theories of gravity
At this point, the variational approach followed above becomes useless for obtaining the corresponding boundary terms because the connection coefficients are fixed at the boundary and, consequently, their variations on the hypersurface $\delta\mathcal{V}$ vanish by definition, $\delta\Gamma=0$. Nevertheless, such boundary terms are present, as shown for instance in Ref. Saez-ChillonGomez:2020afj ; BarberoG:2021cei . A possible way out of this seemingly contradiction is by considering the independent connection as being the Levi-Civita connection of an undetermined metric tensor, which becomes $q_{\mu\nu}$ after solving the field equations for the connection (12). However, such an approach would be against the philosophy of the variational principle, because it would be restricting the possible variations of the connection to those that satisfy the field equations. In fact, the connection has a rank-3 tensorial component whose variations cannot be completely encoded in the variations of a rank-2 tensor. For this reason, a more general approach is necessary to make progress in the understanding of this problem. Next we will proceed by decomposing the Ricci tensor in the $3+1$ formulation. An additional covariant discussion will be provided afterwards.
@@ III $3+1$ decomposition and boundary terms
In order to establish the decomposition of the Ricci tensor and curvature in terms of the properties of a hypersurface (or a family of hypersurfaces), firstly we shall need to decompose the affine connection as:
@@
where $L_{\mu\nu}^{\lambda}$ represents the part of the Levi-Civita connection built with the metric $q_{\mu\nu}$, while the rank-3 tensor $A_{\mu\nu}^{\lambda}$ contains all (so far undefined) extra terms. The choice of $q_{\mu\nu}$ as the metric associated to $L_{\mu\nu}^{\lambda}$ (instead of $g_{\mu\nu}$ or any other metric tensor) is motivated by the resemblance that our effective action (4) has with respect to GR when formulated in terms of it and, obviously, from the fact that the field equations will choose such a connection, as given by Eq.(12). Under this decomposition of the affine connection, the Riemann tensor can be expressed as:
@@
This action is none other than the same action for GR with the GHY term but now defined in terms of the metric tensor $q$ instead of the space-time metric $g$. As shown in the next section, this might imply important and different consequences than in GR.
@@ IV Covariant description of the boundary term revisited
Let us now turn again to Eq.(15) and reconsidered it from the perspective of the splitting (16), such that $\delta\Gamma^{\lambda}_{\mu\nu}=\delta L^{\lambda}_{\mu\nu}+\delta A^{\lambda}_{\mu\nu}$. By construction, the contribution coming from the $q$-dependent $\delta L^{\lambda}_{\mu\nu}$ term is the usual one found in GR for a metric denoted $q_{\mu\nu}$, while the contributions due to the arbitrary connection become
@@
which is exactly the term obtained in the previous section (40) and compensates the variations given in (42). We thus see that, as shown in the previous section by other means, the generalization (46) of the GHY term does indeed compensate the boundary terms generated by the independent connection, both from the metric part and from the non-metric part. Note that the way this occurs is very interesting, as it implies that the $\delta A^{\alpha}_{\mu\nu}$ terms coming from the bulk action and from the GHY term cancel identically. In other words, unlike in Eq. (15), one does not need to impose the vanishing of $\delta A^{\alpha}_{\mu\nu}$ on the boundary to have a well defined variational formulation because with the proposal (46) all such terms vanish identically for arbitrary variations.
@@ VI Conclusions
Along this paper, the Hamiltonian formulation of RBG theories has been obtained. To do so, we have started by using the corresponding mapping proposed in Afonso:2018bpv , by which the space-time metric is replaced in the action by another metric tensor that leads to a linear Lagrangian with respect to the Ricci tensor and a similar form as the usual Hilbert-Einstein action coupling to a modified matter sector. Such a tool has turned out to very useful as it allows to both analyze formal aspects and find explicit solutions of these theories in a much simpler way.
@@
Anchored on this procedure, we built the Hamiltonian formulation first by expressing the Riemann tensor in terms of projections onto a (family of) hypersurface(s), leading to the $3+1$ decomposition of the theory by expressing the Ricci tensor/scalar in terms of the extrinsic and intrinsic curvature of the hypersurface. As a consequence, some new terms might arise in comparison to those of GR, as the presence of a 3-rank tensor in the connection components introduces new terms in such decomposition. However, since the field equations make such tensor null, similar expressions as those of GR are obtained. However, one should note that such decomposition is now performed in terms of a metric tensor that does not coincide with the space-time metric one, but are instead related by a tensorial transformation. The corresponding GHY boundary term is obtained for any RBG theory, which in particular recovers the one obtained previously in the literature for the particular $f(R)$ gravity case Saez-ChillonGomez:2020afj .
@@
Next, by following the standard procedure as in GR, the ADM decomposition as well as the Hamiltonian formulation of the theory were carried out. The Hamiltonian constraints are then obtained, which turn out to coincide with the ones in GR, making the boundary terms the only non-zero contributions to the ADM Hamiltonian. The point is that for the general case, the ADM energy is sourced by the two boundary integrals as given in (72), contrary to the GR case where the second integral cancelled out. Such contributions now depend on the extrinsic curvature as defined in terms of the connection compatible to the metric tensor $q_{\mu\nu}$ and, consequently, it includes contributions that depend on the derivatives of the action through the matrix that relates both metrics via Eq.(3). While such a result were expected, as it already occurs for instance in metric and Palatini $f(R)$ theories with the boundary contributions, it also affects the ADM energy. Indeed, as shown for the simple case of Schwarzschild space-time where the contribution from the second integral is removed, this leads to a contribution on the ADM energy that would vary from one action to other. In addition, the way this might be interpreted can be as follows: the contribution to an effective gravitational constant $G_{eff}=\Omega G$, such that the energy is then given by the mass of the central object. However, this effective gravitational constant will vary from one system to another. On the other hand, the ADM energy (80) might be read off as a sort of self-contribution of gravity to the energy. The latter would lead to some exotic consequences as possible naked singularities keeping the energy of the system positive.

### v2

@@ I Introduction
There is one issue that has remained unsolved over all these years though, namely, the development of a well-defined Hamiltonian formulation for any RBG. The main pitfall behind lied on formal aspects, and mainly on establishing a well-defined variational principle for these theories. The problem has been extensively analysed in GR, where also some issues arise when considering null hypersurfaces Parattu:2015gga , and new techniques have been developed over the last years Chakraborty:2016yna . The point in RBG theories is that as far as one assumes fixed values of the fields at the boundaries of the volume over which the action is integrated, the boundaries terms, as the Gibbons-Hawking-York (GHY) term of GR, are removed, since variations of the connection turn out null at the boundaries. Nevertheless, the presence of such boundary terms have been shown for the case of Palatini $f(R)$ gravities to be fundamental, among other stuff, to recover the correct expression for the Schwarzschild black hole entropy Saez-ChillonGomez:2020afj . In the same way, such boundary terms, analogs of the GHY term, have been obtained by different approaches BarberoG:2021cei ; Gomez:2021roj .
The aim of this paper is to go one step further in the enlightening of this topic with respect to previous analyses. Indeed, here we shall consider a completely general RBG theory and we settle to obtain first the corresponding boundary terms in the gravitational action and, second, to establish a well-defined Hamiltonian formalism for this class of theories. To do so, the mapping of RGB into GR-like form is used in order to transform the general gravitational action into one with just linear terms of the Ricci tensor Afonso:2018bpv . Then, by decomposing the free symmetric connection into two parts, a fiducial connection compatible with a metric plus a rank-3 tensor, the Riemann and Ricci tensors are expressed in terms of these two objects for foliating the corresponding manifold into a family of hypersurfaces. The latter allows to express such tensors in terms of extrinsic and intrinsic magnitudes of the hypersurfaces, the so-called Gauss-Codazzi equations (also known as the $3+1$ decomposition), already implemented successfully in Palatini $f(R)$ gravities Gomez:2021roj . Finally, the Hamiltonian formulation is obtained by following the standard procedure of implementing the Arnowitt-Deser-Misner (ADM) decomposition. Results mimic GR formulation, as expected due to the mapping. However, important different features arise, as shown when computing the ADM energy for the Schwarzschild black hole. Moreover, one should note that the same boundary terms can be obtained by applying variations directly into the action without any mapping to GR. For the latter, the crucial point is to decompose the connection into two parts, one representing a fiducial connection compatible with an unspecified metric tensor, and a second part that consists on a rank-3 tensor whose field equations lead to a trivial result.
@@ II Variational principle in Ricci-based theories of gravity
At this point, the variational approach followed above becomes useless for obtaining the corresponding boundary terms because the connection coefficients are fixed at the boundary and, consequently, their variations on the hypersurface $\delta\mathcal{V}$ vanish by definition, $\delta\Gamma=0$. Nevertheless, such boundary terms are present, as shown for instance in Ref. Saez-ChillonGomez:2020afj ; BarberoG:2021cei . A possible way out of this seemingly contradiction is by considering the independent connection as being the Levi-Civita connection of an undetermined metric tensor, which becomes $q_{\mu\nu}$ after solving the field equations for the connection (12). However, such an approach would be against the philosophy of the variational principle, because it would be restricting the possible variations of the connection to those that satisfy the field equations. In fact, the variations of the connection can be represented by a rank-3 tensorial component whose variations cannot be completely encoded in the variations of a rank-2 tensor. For this reason, a more general approach is necessary to make progress in the understanding of this problem. Next we will proceed by decomposing the Ricci tensor in the $3+1$ formulation. An additional covariant discussion will be provided afterwards.
@@ III $3+1$ decomposition and boundary terms
In order to establish the decomposition of the Ricci tensor and curvature in terms of the properties of a hypersurface (or a family of hypersurfaces), firstly we shall decompose the affine connection as
@@
where $L_{\mu\nu}^{\lambda}$ represents a fiducial Levi-Civita connection built with the metric $q_{\mu\nu}$, while the rank-3 tensor $A_{\mu\nu}^{\lambda}$ contains all (so far undefined) extra terms. The choice of $q_{\mu\nu}$ as the metric associated to $L_{\mu\nu}^{\lambda}$ (instead of $g_{\mu\nu}$ or any other metric tensor) is motivated by the resemblance that our effective action (4) has with respect to GR when formulated in terms of it and, obviously, from the fact that the field equations will choose such a connection, as given by Eq.(12). Note that the decomposition (16) puts forward that when a fiducial connection is defined, such as $L_{\mu\nu}^{\lambda}$ here, there exists a bijection between the difference of two connections and the space of rank-3 tensors. The space of connections is an affine space, while the space of tensors is a vector space and, generally, it is easier and clearer to work with vector spaces. That is why we have followed this approach and, as we will see, it will proof useful in this discussion. Under this decomposition of the affine connection, the Riemann tensor can be expressed as:
@@
This action is none other than the same action for GR with the GHY term but now defined in terms of the metric tensor $q$ instead of the space-time metric $g$. As shown in the next section, this might imply important and different consequences as compared to GR.
@@ IV Covariant description of the boundary term revisited
Let us now turn again to Eq.(15) and reconsider it from the perspective of the splitting (16), such that the variation $\delta\Gamma^{\lambda}_{\mu\nu}$ can be split as the sum of the variation of the fiducial connection $L^{\lambda}_{\mu\nu}$ plus the variation of the rank-3 tensor $A^{\lambda}_{\mu\nu}$, namely, $\delta\Gamma^{\lambda}_{\mu\nu}=\delta L^{\lambda}_{\mu\nu}+\delta A^{\lambda}_{\mu\nu}$. By construction, the contribution coming from the $q$-dependent $\delta L^{\lambda}_{\mu\nu}$ term is the usual one found in GR for a metric denoted $q_{\mu\nu}$, while the contributions due to the arbitrary connection become
@@
which is exactly the term obtained in the previous section (40) and compensates the variations given in (42). We thus see that, as shown in the previous section by other means, the generalization (46) of the GHY term does indeed compensate the boundary terms generated by the independent connection, both from the metric part and from the non-metric part. Note that the way this occurs is very interesting, as it implies that the $\delta A^{\alpha}_{\mu\nu}$ terms coming from the bulk action and from the GHY term cancel identically. In other words, unlike in Eq. (15), one does not need to impose the vanishing of $\delta A^{\alpha}_{\mu\nu}$ on the boundary to have a well defined variational formulation because with the proposal (46) all such terms vanish identically for arbitrary variations. Note also that the same results are obtained even in absence of any mapping of the gravitational action into a GR-like action as far as one splits the independent connection as (16) and then, takes variations over the metric tensor $q$ and the independent 3-rank tensor $A$.
@@ VI Conclusions
In this paper, the Hamiltonian formulation of RBG theories has been obtained together with an analysis of its boundary terms. To do so, we have started by using the corresponding mapping proposed in Afonso:2018bpv , by which the space-time metric is replaced in the action by another metric tensor that leads to a linear Lagrangian with respect to the Ricci tensor and a similar form as the usual Hilbert-Einstein action coupling to a modified matter sector. Such a tool has turned out to very useful as it allows to both analyse formal aspects and find explicit solutions of these theories in a much simpler way.
The decomposition of the connection $\Gamma_{\mu\nu}^{\lambda}$ given in Eq. (16) allowed us to interpret its variation as the variation of the fiducial connection $L_{\mu\nu}^{\lambda}$ plus the variation of the rank-3 tensor $A_{\mu\nu}^{\lambda}$. This puts forward that the variation of a connection cannot be regarded as just the variation of an ordinary tensorial field, because in that case one would miss the additional contribution that we found in $L_{\mu\nu}^{\lambda}$, which is crucial to obtain the correct boundary terms. This clarifies the apparent inconsistency pointed out at the end of section II, when the variation was implemented assuming that $\delta\Gamma_{\mu\nu}^{\lambda}\sim\delta A_{\mu\nu}^{\lambda}$.
@@
Anchored on this procedure, we built the Hamiltonian formulation first by expressing the Riemann tensor in terms of projections onto a (family of) hypersurface(s), leading to the $3+1$ decomposition of the theory by expressing the Ricci tensor/scalar in terms of the extrinsic and intrinsic curvature of the hypersurface. As a consequence, some new terms might arise in comparison to those of GR, as the presence of a rank-3 tensor in the connection components introduces new terms in such decomposition. However, since the field equations make such tensor null, similar expressions as those of GR are obtained. However, one should note that such decomposition is now performed in terms of a metric tensor that does not coincide with the space-time metric, but are instead related by a tensorial transformation. The corresponding GHY boundary term is obtained for any RBG theory, which in particular recovers the one obtained previously in the literature for the particular $f(R)$ gravity case Saez-ChillonGomez:2020afj .
@@
Next, by following the standard procedure as in GR, the ADM decomposition as well as the Hamiltonian formulation of the theory were carried out. The Hamiltonian constraints are then obtained, which turn out to coincide with the ones in GR, making the boundary terms the only non-zero contributions to the ADM Hamiltonian. The point is that for the general case, the ADM energy is sourced by the two boundary integrals as given in (72), contrary to the GR case where the second integral cancelled out. Such contributions now depend on the extrinsic curvature as defined in terms of the connection compatible to the metric tensor $q_{\mu\nu}$ and, consequently, it includes contributions that depend on the derivatives of the action through the (matter-dependent) matrix that relates both metrics via Eq.(3). While such a result was expected, as it already occurs for instance in metric and Palatini $f(R)$ theories with the boundary contributions, it also affects the ADM energy. Indeed, as shown for the simple case of Schwarzschild space-time where the contribution from the second integral is removed, this leads to a contribution on the ADM energy that would vary from one action to other. In addition, the way this might be interpreted can be as follows: the contribution to an effective gravitational constant $G_{eff}=\Omega G$, such that the energy is then given by the mass of the central object. However, this effective gravitational constant will vary from one system to another. On the other hand, the ADM energy (80) might be read off as a sort of self-contribution of gravity to the energy. The latter would lead to some exotic consequences as possible naked singularities keeping the energy of the system positive.